​ An investment has an installed cost of $673.658

An investment has an installed cost of $673.658. The cash flows over the four year life of the investment are projected to be $228,701, $281,182, $219,209, and $190,376. 


Requirement 1: 

If the discount rate is zero, what is the NPV? (Do not round intermediate calculations.) 


Requirement 2: 

If the discount rate is infinite, what is the NPV? (Do not round intermediate calculations. Negative amount should be indicated by a minus sign.) 


Requirement 3: 

At what discount rate is the NPV just equal to zero? (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g. 32.16).) 
Answer #1

Requirement 1: NPV at discount rate of 0%
NPV=-Initial cost + Present value of future cash flows
Present value of future cash flows=(Cash flow in year 1)/(1+Discount rate)^1+(Cash flow in year 2)/(1+Discount rate)^2+(Cash flow in year 3)/(1+Discount rate)^3+
(Cash flow in year 4)/(1+Discount rate)^4

=-673658+228701/(1+0)^1+281182/(1+0)^2+219209/(1+0)^3+190376/(1+0)^4
=-673658+919468
=245810

Requirement 2: NPV at discount rate of infinity
=-673658+228701/(1+Infinity)^1+281182/(1+Infinity)^2+219209/(1+Infinity)^3+190376/(1+Infinity)^4
=-673658+0+0+0+0
=-673658

Requirement 3:
The value of NPV becomes zero when the discount rate is equal to internal rate of return (or IRR)


So, when the discount rate is 14.30%, the value of NPV will be zero